# When Location Shapes Choice:

### v3

@@ 4.3 Tightness of Analysis
The next theorem proves the tightness of analysis of our approximation algorithms for the Placement problem.
@@
To prove the above theorem, we will provide three technical lemmas. We include the proofs of these lemmas in Appendix 9.1.
Lemma 4.13
For each $k\in[m]$, the optimal assortment $S^{*}_{k}$ of size at most $k$ in our instance is the set of the $k$ products with preference weights $v_{k}$.
Lemma 4.14
Let $k\in[m]$. Consider $X_{k}$ a placement of the products in $S^{*}_{k}$ to the locations of the line (possibly a randomized placement). Let $\mathcal{R}(X_{k})$ be the revenue obtained by this placement. Then, $\mathcal{R}(X_{k})=O(1)$.
Lemma 4.15
Let $X$ be a placement that, for each $j\in[m]$, assigns the product of preference weight $u_{j}$ to location $j$. Then, $\mathcal{R}(X)=\Omega(\log^{1-\epsilon}m)$.
Combining the results of Lemma 4.14 and 4.15 directly yields Theorem 4.12.
@@ 4.4 Deterministic Algorithm for Markov Choice Model
For every $k\in[m]$, there is a $(1-1/e)$-approximation algorithm for finding the optimal placement on the reduced ground set $S^{*}_{k}$ when $\phi$ is a Markov choice model.
We include the proof of Lemma 4.16 in Appendix 9.2. Let $\hat{X}_{k}$ denote a $(1-1/e)$-approximate placement for $S^{*}_{k}$. Consider the placement solution,
@@ 5 Numerical Experiments
In this section, we describe numerical experiments on synthetic data to evaluate the performance of our algorithms for the Placement problem.
@@ 5.1 Experimental Setup
This setup is inspired by the layout of a physical store, where the corners represent entrances and exits, and customers navigate through the store via random walks. We approximate the browsing distribution by generating $\Gamma$ random walks. For our numerical experiments, we fix $n=20$ and $\Gamma=10$. We vary $d\in\{1,2,3,4,5\}$ which means that $m\in\{4,9,16,25,36\}$.
@@ 5.2 Results
For each value of $n$ and $m$, we sample 100 instances of the Placement problem according to the distributions described above. For each instance, we solve the Placement problem optimally by developing a mixed integer linear programming (MILP) formulation. The details of our formulation are provided in Appendix 11. Note that the MILP formulation uses big-M constrains and can accommodate general browsing distributions. The formulation can be solved efficiently for relatively small instances but it takes a significant amount of time to solve larger ones. We compare the optimal solution with the solution generated by our algorithms. For the line, we use the deterministic algorithm described in Section 4.1. For the two-dimensional grid, we use the general randomized algorithm described in Section 4.2.
Placement on a Line: The results are reported in Table 1. The first and second columns in Table 1 present the values of $m$ and $n$. The third and fourth columns provide the ratio of the objective value obtained by our algorithm to the optimal objective value. Out of the 100 instances, we report the minimum ratio and the mean ratio. The fifth column shows the running time of our algorithm in seconds, while the sixth column reports the running time of the MILP formulation, also in seconds.
@@
$n$ $m$ Ratio Time ALG Time OPT
Min Mean
@@
Table 1: Performance of our algorithm on a Line
$n$ $d$ $m$ Ratio Time ALG Time OPT % Instances $<$ 1H
Min Mean
10 1 4 0.812 0.941 0.05 0.23 100
2 9 0.866 0.938 0.11 3.59 100
3 16 0.872 0.934 0.22 7.98 100
4 25 0.884 0.936 0.35 3.05 100
5 36 0.894 0.944 0.52 1.25 100
15 1 4 0.833 0.951 0.05 0.69 100
2 9 0.892 0.952 0.13 220.85 99
3 16 0.899 0.947 0.22 415.83 91
4 25 0.892 0.942 0.35 142.51 95
5 36 0.898 0.944 0.53 79.51 100
20 1 4 0.840 0.958 0.05 2.08 100
2 9 0.905 0.957 0.14 576.01 80
3 16 0.929 0.951 0.25 875.19 30
4 25 0.916 0.949 0.38 722.37 52
5 36 0.904 0.948 0.59 519.23 70
Table 2: Performance of our algorithm on a two-dimensional grid
@@
Placement on a Two-Dimensional Grid: The results are reported in Table 2. The first, second, and third columns in Table 2 present the values of $n$, $d$, and $m$. For each pair $(n,d)$, we run 100 instances of the placement problem. We observe that the running time of the MILP used to compute the optimal solution of the placement problem can be extremely high; therefore, we limit the running time to 1 hour. In the last column of the table, we report the percentage of instances that are solved to optimality within the 1-hour limit. For each instance where the optimal solution is computed within the time limit, we generate a solution using the algorithm described in Section 4.2, and then calculate the minimum and average ratios of the objective values over all such instances. Note that we execute our randomized algorithm once for each instance. The third and fourth columns of the table report these ratios: the minimum ratio and the mean ratio of the objective value obtained by our algorithm to the optimal objective value, evaluated over the instances solved to optimality. The fifth column shows the average running time of our algorithm in seconds. The sixth column reports the average running time of the MILP approach, also in seconds, with the average calculated only over the instances that are solved to optimality.
@@
Our algorithm achieves an average ratio of 0.94 across all values of $n$ and $d$. In the worst case, the minimum ratio is 0.81, achieved for the smallest instance with $n=10$ and $d=1$. In fact, as we increase the problem size, it seems that the performance of the algorithm improves. Note that the ratios are slightly smaller than those for Placement on a Line, but they are still very high. In theory, our algorithm achieves a worst-case ratio of $(1-1/e)/\log(m+1)$ for the grid (as shown in Theorem 4.6), but we observe significantly better empirical performance. Furthermore, our algorithm is extremely fast, with a running time of less than one second. In contrast, the running time of the MILP approach increases exponentially with $m$ and $n$, becoming significantly higher as the dimensions increase. In fact, obtaining the optimal solution becomes extremely challenging as the dimensions grow. For example, while we are able to solve all instances for $n=10$ and most instances for $n=15$ within the 1-hour time limit, solving instances for $n=20$ becomes much harder. For instance, when $n=20$ and $d=3$, only 30 instances converge to optimality within the 1 hour time limit. The MILP formulation, with its large number of variables and constraints and big-$M$ constraints, becomes impractical. This highlights the necessity of our algorithm, which can quickly compute good solutions.
@@ 6 Conclusion
We study the problem of finding an optimal placement of substitutable products by introducing a model where customer choice is a two stage process. First, the customer browses the store and visits a subset of display locations according to a known browsing distribution. Then, the customer chooses from the assortment of products located at the visited locations according to a discrete choice model. We consider the problem for a general browsing distribution and general discrete choice model, a setting for which no approximation results were known in the literature (to the best of our knowledge). We leverage the existing results for cardinality constrained assortment optimization to obtain provably good placement of products in this general setting. In particular, we develop an approximation algorithm that uses an $\alpha$-approximation algorithm for cardinality constrained assortment optimization to generate a $\frac{\Theta(\alpha)}{\log_{2}\rho}$-approximation for the placement problem for any browsing distribution. Our algorithm is randomized and we show that it can be derandomized for the Markov choice model. When products are homogeneously priced, we show that the problem is an instance of monotone submodular function maximization subject to a partition matroid constraint. This gives a $(1-1/e)$-approximation for our setting in the special case of homogeneous prices. We show that unless P=NP, no polynomial time algorithm can beat this guarantee.
Improving the $\frac{1}{\log_{2}\rho}$ factor in our guarantee or showing that no efficient algorithm has a better guarantee remains a challenging open problem. Our model assumes that the browsing distribution does not depend on the placement solution. This may not be true in general, especially in settings such as cultural institutions (Aouad et al. (2022)), and finding an approximation algorithm without this assumption remains a challenging open problem.

### v4

@@ 4.3 Tightness of Analysis
The next theorem proves the tightness of analysis of our approximation algorithms for the Placement problem. The proof of Theorem 4.11 is given in Appendix 9.
@@ 4.4 Deterministic Algorithm for Markov Choice Model
For every $k\in[m]$, there is a (11/e)(1-1/e)-approximation algorithm for finding the optimal placement on the reduced ground set $S^{*}_{k}$ when $\phi$ is a Markov choice model.
We include the proof of Lemma 4.12 in Appendix 9.6. Let $\hat{X}_{k}$ denote a (11/e)(1-1/e)-approximate placement for $S^{*}_{k}$. Consider the placement solution,
@@ 5 Numerical Experiments
In this section, we present numerical experiments on synthetic data to evaluate the performance of our algorithm for the Placement problem (Sections 5.2.1 and 5.2.2), as well as to assess the impact of product repetition (Section 5.2.3).
@@ 5.1 Experimental Setup
This setup is inspired by the layout of a physical store, where the corners represent entrances and exits, and customers navigate through the store via random walks. We approximate the browsing distribution by generating $\Gamma$ random walks, each with equal probability. For our numerical experiments, we choose $n\in\{10,15,20\}$ and $\Gamma=10$. We vary $d\in\{1,2,3,4,5\}$ which means that $m\in\{4,9,16,25,36\}$.
@@ 5.2 Results
For each value of $n$ and $m$, we sample 100 instances of the Placement problem according to the distributions described above. For each instance, we solve the Placement problem optimally by developing a mixed integer linear programming (MILP) formulation. The details of our formulation are provided in Appendix 11. Note that the MILP formulation uses big-M constraints and can accommodate general browsing distributions. The formulation can be solved efficiently for relatively small instances but it takes a significant amount of time to solve larger ones. Note that because the browsing distribution is approximated via a finite set of sampled random walks, the MILP we solve is in fact an instance of a sample average approximation (SAA)–based integer program, i.e., the browsing distribution is represented by the empirical distribution over the sampled paths. We compare the optimal solution with the solution generated by our algorithms. For the line, we use the deterministic algorithm described in Section 4.1. For the two-dimensional grid, we use the general randomized algorithm described in Section 4.2. We use Gurobi 10.0.2 as the integer programming solver, and the numerical experiments were conducted in Python on a MacBook Pro equipped with a 2.3-GHz 8-core Intel Core i9 processor and 32 GB RAM.
5.2.1 Placement on a Line.
The results are reported in Table 1. The first and second columns in Table 1 present the values of $m$ and $n$. The third and fourth columns provide the ratio of the objective value obtained by our algorithm to the optimal objective value. Out of the 100 instances, we report the minimum ratio and the mean ratio. The fifth column shows the running time of our algorithm in seconds, while the sixth column reports the running time of the MILP formulation, also in seconds.
@@
$n$ $m$ Ratio Time(s)
Min Mean ALG OPT
@@
Table 1: Results of experiments on Line graphs. For each $(n,m)$ pair, we report the minimum and mean performance of ALG relative to OPT and the average running times.
$n$ $d$ $m$ Ratio Time(s) % Instances Occurrences Repet. Gain (%)
Min Mean ALG OPT <<1H ALG OPT Mean Max
10 1 4 0.740 0.960 0.12 0.12 100 2.01 2.07 21.1 77.8
2 9 0.858 0.958 0.14 1.40 100 3.93 2.94 31.5 63.1
3 16 0.839 0.962 0.15 2.42 100 6.42 4.10 36.2 71.3
4 25 0.882 0.968 0.16 2.02 100 9.86 5.42 42.0 129.7
5 36 0.877 0.968 0.16 28.47 100 15.7 6.95 43.6 81.7
15 1 4 0.841 0.960 0.19 0.25 100 1.88 1.96 18.1 39.8
2 9 0.868 0.958 0.21 11.79 100 3.47 2.72 27.6 74.5
3 16 0.897 0.958 0.22 18.08 100 5.55 3.84 33.2 65.3
4 25 0.844 0.965 0.26 37.63 100 8.64 5.39 35.2 81.4
5 36 0.855 0.964 0.26 27.05 100 13.3 6.80 37.6 79.9
20 1 4 0.851 0.960 0.27 0.35 100 1.67 1.96 14.1 50.8
2 9 0.877 0.953 0.30 24.27 100 3.29 2.61 24.4 49.1
3 16 0.861 0.959 0.32 165.45 94 5.70 3.72 30.5 51.6
4 25 0.897 0.967 0.33 134.46 89 8.41 4.91 30.7 49.0
5 36 0.882 0.967 0.39 225.64 89 11.9 6.45 35.5 56.5
Table 2: Results of experiments on two-dimensional grids. For each $(n,d)$ pair, we report the minimum and mean performance of ALG relative to OPT, average running times, and the percentage of instances solved by the MILP within one hour. The last few columns show the extent of product repetition and the revenue gains from allowing repetition (see Section 5.2.3).
5.2.2 Placement on a Two-Dimensional Grid.
@@
The results are reported in Table 2. The first, second, and third columns in Table 2 present the values of $n$, $d$, and $m$. For each pair $(n,d)$, we run 100 instances of the placement problem. We observe that the running time of the MILP used to compute the optimal solution of the placement problem can be extremely high; therefore, we limit the running time to 1 hour. In the eight column of the table (% Instances <<1H), we report the percentage of instances that are solved to optimality within the 1-hour limit. For each instance where the optimal solution is computed within the time limit, we generate a solution using the algorithm described in Section 4.2, and then calculate the minimum and average ratios of the objective values over all such instances. Note that we execute our randomized algorithm once for each instance. The fourth and fifth columns of the table (Ratio Min and Ratio Mean) report these ratios: the minimum ratio and the mean ratio of the objective value obtained by our algorithm to the optimal objective value, evaluated over the instances solved to optimality. The sixth column (Time(s)
Algorithm 1
) shows the average running time of our algorithm in seconds. The seventh column (Time(s) OPT) reports the average running time of the MILP approach, also in seconds, with the average calculated only over the instances that are solved to optimality.
@@
Our algorithm achieves an average ratio of 0.96 across all values of $n$ and $d$. In the worst case, the minimum ratio is 0.74, achieved for the smallest instance with $n=10$ and $d=1$. In fact, as we increase the problem size, it seems that the performance of the algorithm improves. Note that the ratios are slightly smaller than those for Placement on a Line, but they are still very high. In theory, our algorithm achieves a worst-case ratio of $(1-1/e)/\log(m+1)$ for the grid (as shown in Theorem 4.7), but we observe significantly better empirical performance. Furthermore, our algorithm is extremely fast, with a running time of less than one second. In contrast, the running time of the MILP approach increases exponentially with $m$ and $n$, becoming significantly higher as the dimensions increase. In fact, obtaining the optimal solution becomes extremely challenging as the dimensions grow. For example, while we are able to solve all instances for $n=10$ and $n=15$ within the 1-hour time limit, solving instances for $n=20$ becomes much harder. For instance, when $n=20$ and $d=5$, 11 instances did not converge to optimality within the 1 hour time limit. The MILP formulation, with its large number of variables and constraints and big-$M$ constraints, becomes impractical. This highlights the necessity of our algorithm, which can quickly compute good solutions.
5.2.3 Value of Repetition in the Grid.
Recall that in our Placement problem we allow products to be repeated in multiple locations. In the Placement on a Line setting, such repetition does not provide any benefit (see Remark 2.4). In the worst case, not allowing product repetition can lead to significantly lower revenue (see Example 4.6). To quantify the significance of repetition beyond the worst case, we numerically evaluate its effect in the grid setting.
We first calculate for each product the number of times it is duplicated across locations (we refer to this as its number of occurrences), and we average this quantity over products with at least one occurrence. We compute this measure both for the optimal solution (OPT) and for the solution returned by our algorithm (ALG) and report these averages in the ninth and tenth columns of Table 2 (Occurrences
Algorithm 2
and Occurrences OPT). The results show that both OPT and ALG exhibit duplication, with an average number of occurrences equal to 4.12 for OPT and 6.77 for ALG across all our instances. As expected, when the number of locations increases while the number of products is fixed, we observe more duplication; conversely, increasing the number of products for a fixed number of locations reduces the extent of duplication. Furthermore, OPT systematically uses fewer duplicates than ALG. This is natural: OPT leverages global structural considerations to minimize unnecessary repetition, while ALG, by construction, samples uniformly from a candidate subset at each location, which tends to generate more duplicates.
To evaluate the value of repetition, we compare the optimal revenue from the Placement problem (i.e., the optimal objective value of Placement obtained through the MILP in Appendix E) with the optimal revenue of the same problem when repetition is disallowed. The latter is formulated by simply adding a constraint to the MILP in Appendix E requiring that each product appear in at most one location; we denote its optimal objective value by $\text{OPT}_{\text{norep}}$. By construction, $\text{OPT}\geq\text{OPT}_{\text{norep}}$. We define the Repetition Gain as $\frac{\text{OPT}-\text{OPT}_{\text{norep}}}{\text{OPT}_{\text{norep}}},$ and report this quantity in the last two columns of Table 2 (Repet. Gain (%)), both as an average over all 100 instances (Mean) and as the maximum observed value (Max). The results indicate that repetition provides a significant benefit. On average, the gain ranges from roughly 14% to 43%, depending on the values of $n$ and $m$, and in some cases the maximum gain can be much higher. For example, when $n=10$ and $d=4$, the repetition gain reaches 129% in average, representing the highest value observed in our experiments. These findings confirm that, unlike in the line setting where duplicates play no role, repetition in the grid setting can substantially improve revenues by amplifying the placement of revenue generating products across multiple visible locations.
Structure of Solutions: To further understand the structure of the solutions, we look at how product prices influence the locations where products are placed and the number of occurrences. We report our results in Figures 6 and 7 in Appendix 12, and summarize our main findings here. Our experiments reveal several consistent patterns in the structure of OPT and
Algorithm 3
. Both exhibit a clear “revenue-ordered” structure: the highest-revenue products are duplicated most often and are preferentially assigned to the most visible locations, with duplication frequency and visibility decreasing as product revenue decreases. While this ordering is stable across different values of $n$ and $m$, ALG tends to employ more duplication than OPT, reflecting its randomized design. Overall, these findings confirm that both methods exploit the MNL model’s tendency to favor revenue-ordered assortments, with ALG trading off optimality for simplicity by duplicating high-revenue products more aggressively.
@@ 6 Conclusion
We study the problem of finding an optimal placement of substitutable products by introducing a model where customer choice is a two stage process. First, the customer browses the store and visits a subset of display locations according to a known browsing distribution. Then, the customer chooses from the assortment of products located at the visited locations according to a discrete choice model. We consider the problem for a general browsing distribution and general discrete choice model, a setting for which no approximation results were known in the literature (to the best of our knowledge). We leverage the existing results for cardinality constrained assortment optimization to obtain provably good placement of products in this general setting. In particular, we develop an approximation algorithm that uses an $\alpha$-approximation algorithm for cardinality constrained assortment optimization to generate a $\frac{\Theta(\alpha)}{\log_{2}\rho}$-approximation for the placement problem for any browsing distribution. Our algorithm is randomized and we show that it can be derandomized for the Markov choice model. When products are homogeneously priced, we show that the problem is an instance of monotone submodular function maximization subject to a partition matroid constraint. This gives a (11/e)(1-1/e)-approximation for our setting in the special case of homogeneous prices. We show that unless P=NP, no polynomial time algorithm can beat this guarantee.
Improving the $\frac{1}{\log_{2}\rho}$ factor in our guarantee or showing that no efficient algorithm has a better guarantee remains a challenging open problem. Our model assumes that the browsing distribution does not depend on the placement solution. This may not be true in general, especially in settings such as cultural institutions (Aouad et al. (2025a)), and finding an approximation algorithm without this assumption remains a challenging open problem.